OpenStudy (anonymous):

Determine the position of all maximum points using the second-derivative test of y=(x^2/6)-2x+8

OpenStudy (kropot72):

Is this it? \[y=\frac{x ^{2}}{6}-2x+8\]

OpenStudy (anonymous):

first differentiate y w.r.t x and find the extremities. then find 2nd derivative of y. check whether it is -ve . if it is -ve then the extremum represents maximum

OpenStudy (kropot72):

\[f'(x)=\frac{x}{3}-2\] x/3 - 2 = 0 x = 6 \[f''(x)=\frac{1}{3}\] f''(x) is positive therefore x = 6 is at a minimum point. There are no maximum points.